OpenStudy (anonymous):

5x - 1 = x + 15 SHOW WORK PLEASE

OpenStudy (anonymous):

5x-1=x+15 -x -x 4x-1=15 +1 +1 4x=16 /4 /4 X=4

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (anonymous):

you are welcome!!!:)